Suppose a condom is 85% safe. Tony came home with a 36 piece box of condoms to please his girlfriend. Suppose he uses it all, What are the chances of Tony getting a child.
Answer #1

Binomial Distribution

n = Total number of trials = 36

p = Probability of safe in a single trial = 0.85

q = Probability of unsafe in a single trial = 1 - 0.85 = 0.15

Probability that Tony gets a child = P(At least one trial is unsafe) = 1 - P(All trials safe) = 1 - P(X=36)

= 1 - pn = 1 - 0.8536 = 1 - 0.002878 = 0.997122

So,

Answer is:

0.997122